OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)?

OpenStudy (rocklionmba):

what are you trying to find?

OpenStudy (anonymous):

4x - y = 1 4x + y = 1 x - 4y = -13 4x - y = -7

OpenStudy (rocklionmba):

ok. so what your going to hae to do is find the slope. do you know the slope formula?

OpenStudy (anonymous):

not really im bad at math

OpenStudy (anonymous):

@trindawg

OpenStudy (rocklionmba):

the formula is\[\frac{ y_{2}+y_{1} }{x_{2}+x_{1}}\]

OpenStudy (anonymous):

oh so i think its b

OpenStudy (anonymous):

u there

OpenStudy (rocklionmba):

yeah

OpenStudy (anonymous):

so what do u think

OpenStudy (rocklionmba):

let me do some calculations

OpenStudy (rocklionmba):

I don't think so, did you do the formula?

OpenStudy (anonymous):

i tried it

OpenStudy (rocklionmba):

and you got...

OpenStudy (anonymous):

i think b or c

OpenStudy (rocklionmba):

no what did the formula give you

OpenStudy (anonymous):

i really didnt get it

OpenStudy (rocklionmba):

ok. so what you have to do is enter the numbers to get it. if you do the formula correctly, youll get 4 as your slope. now you need to find which of the equations have a slope of 4.

OpenStudy (anonymous):

could you help me find it i only have 4 min left on my test

OpenStudy (rocklionmba):

ok. let me get the answer

OpenStudy (rocklionmba):

umm...

OpenStudy (rocklionmba):

ill just tell you because im busy its probably c